OpenStudy (anonymous):

I'm not sure how to do the precise problems! 11ft;122in and can someone explain how to solve these types of problems please

OpenStudy (phi):

what is the exact problem? 11 feet; 122 inches does not mean much. do you want to compare them, and say which is bigger? if so, you need to use the *same units* (there is a saying, "compare apples to apples") you can change feet to inches by multiplying, like this: \[ 11 \ ft \cdot \frac{ 12 \ in}{1 \ ft} = 11\cdot 12 \ in = 132 \ in\] now you can compare 132 inches to 122 inches